OpenStudy (anonymous):

Find the real or imaginary solutions of each equation by factoring. 81^3-192=0

OpenStudy (anonymous):

@hartnn how do I get rid of the 81?

OpenStudy (solomonzelman):

I messed up... it's not clear what the problem actually is -:( Is it \[81x^3=192~~~?\]

OpenStudy (solomonzelman):

\[81x^3-192=0~~~~~~~~~~~~~~~\frak rig ht~?\]

OpenStudy (anonymous):

It's the second one you posted.

OpenStudy (solomonzelman):

Well, they are the same:) the first one ---> is when I subtract 192 from both sides in the second one. Anyway, \[81x^3-192=0\]add 192 to both sides\[81x^3=192\]\[3 \times 27x^3=3 \times 64\]\[27x^3=64\]\[3^{3}x^3=4^{3}\]\[(3x)^3=4^{3}\]can you finish the problem from here ?

OpenStudy (solomonzelman):

the first phrase, "add" not "subtract" ....

OpenStudy (anonymous):

Where did you get 64?

OpenStudy (solomonzelman):

In my process I re-wrote 192 as 64×3 ( and rewrote 81, as 27×3)

OpenStudy (solomonzelman):

I didn't get that I got\[27x^{3}\]and then re-wrote 27 as 3³